# The in-context inductive biases of vision-language models differ across modalities

### v1

@@ 1 Introduction
We focus on a well-studied inductive bias: the difference between color and shape features in category learning. Many studies show that humans prefer to generalize along shape rather than dimensions like color or texture, from a very young age (Bornstein, 1985; Landau et al., 1988). That is, in ambiguous situations where either shape or color could be the feature determining a category, humans will generally assume that shape is the correct feature to generalize. This tendency in humans has been contrasted with that in convolutional vision models, which tend to prefer lower-level features like texture or color over shape (Geirhos et al., 2019; 2020; Hermann & Lampinen, 2020). Recent studies on more modern vision-language models have found that they have recovered some human-like shape bias (Gavrikov et al., 2024)—and some steerability from language. However, these studies examine the generalization of these models from their training data; not their generalization from categories learned in context. Here, we instead study the shape-color-bias from in-context examples. We also compare to model behavior when the equivalent stimuli are presented described in text rather than directly provided to the model as an image.
@@ 3 Results
Figure 2: Across three category-generalization paradigms, VLMs generalize more by shapes than colors overall (bars above the midline). This bias tends to be amplified when the categories are presented as images (blue), compared to when they are presented in text (orange). However, this pattern is flipped for the odd-one-out task for some models.
Figure 3: The order in which the features are presented in text shifts the VLMs generalization biases; models show some generalization preference toward the feature that is mentioned first (green bars are higher than yellow bars). For one-category tasks, Claude & GPT refuse all generalizations.
Overall differences between images and text: In Fig. 2 we show an overview of the differences between the inductive biases of the models when learning concepts through image or text modality, across the three different paradigms. The first pattern to note is simply that there is a difference between the modalities; that is, the models’ inductive biases seem to differ between the modalities. The overall pattern is for models to exhibit a stronger shape bias when concepts are presented visually than when they are presented in text (i.e., the blue bars are generally higher on the shape bias axis than the orange bars). This is true for all models for both the explicit category-learning tasks, and for the Gemini model for the odd-one-out task. However, the other two models show a surprising inversion of the pattern for the odd-one-out task, with stronger shape biases from text than images.
Feature order in text affects generalization: In Fig. 3 we show that the order in which features are presented in text has a strong effect on how the models generalize. In general, the models are more shape-biased when shape is mentioned first, and less shape-biased if color is mentioned first. Thus, the models seem to be somewhat biased towards generalizing in accordance with the feature that is mentioned first. However, there is still a weak overall tendency to be shape biased on average.
@@ Acknowledgements
We thank Mike Mozer for helpful comments and suggestions.
@@ B.0.1 Single category
We show the results in Fig. 5. Overall, the patterns of behavior are very noisy and inconsistent, though there is an overall tendency to favor shape matches. However, there is a hint of the expected pattern of generalization when comparing the varied-color and varied shape panels; the model tends to generalize more to matching shapes when the color of examples varies (irrespective of modality), but shows a greater degree of color generalization when the colors are consistent. However, the results when both features vary are surprisingly narrow.

### v2

@@ 1 Introduction
We focus on a well-studied inductive bias: the difference between color and shape features in category learning. Many studies show that humans prefer to generalize along shape rather than dimensions like color or texture, from a very young age (Bornstein, 1985; Landau et al., 1988). That is, in ambiguous situations where either shape or color could be the feature determining a category, humans will generally assume that shape is the correct feature to generalize. This tendency in humans has been contrasted with that in convolutional vision models, which tend to prefer lower-level features like texture or color over shape (Geirhos et al., 2019; 2020; Hermann & Lampinen, 2020).
Recent studies on more modern vision-language models have found that they have recovered some human-like shape bias (Gavrikov et al., 2024)—and some steerability from language. Another recent work (Verhoef et al., 2024) found that vision language models do not show the phonological biases in mapping shapes to labels that humans do. However, these studies examine the zero-shot generalization of these models from their training data; not their generalization from categories learned in context. Here, we instead study the shape-color-bias from in-context examples. We also compare to model behavior when the equivalent stimuli are presented described in text rather than directly provided to the model as an image.
@@ 3 Results
Figure 2: Across three category-generalization paradigms, VLMs generalize more by shapes than colors overall (bars above the midline). This bias tends to be amplified when the categories are presented as images (blue), compared to when they are presented in text (orange). However, this pattern is flipped for the odd-one-out task for some models. (Errorbars are bootstrapped 95%-CIs.)
Figure 3: The order in which the features are presented in text shifts the VLMs generalization biases; models show some generalization preference toward the feature that is mentioned first (green bars are higher than yellow bars). For one-category tasks, Claude & GPT refuse all generalizations; hence they show no text-order differences. (Errorbars are boostrapped 95%-CIs.)
Overall differences between images and text: In Fig. 2 we show an overview of the differences between the inductive biases of the models when learning concepts through image or text modality, across the three different paradigms. The first pattern to note is simply that there is a difference between the modalities; that is, the models’ inductive biases seem to differ between the modalities. This overall effect is statistically significant in a regression controlling for task type in all models (respectively; $t(823)=12.478,p<0.001$; $t(823)=2.432,p=0.015$; $t(823)=8.837,p<0.001$). The overall pattern is for models to exhibit a stronger shape bias when concepts are presented visually than when they are presented in text (i.e., the blue bars are generally higher on the shape bias axis than the orange bars). This is true for all models for both the explicit category-learning tasks, and for the Gemini model for the odd-one-out task. However, the other two models show a surprising inversion of the pattern for the odd-one-out task, with stronger shape biases from text than images (though the interaction is only statistically significant for Claude, $t(821)=2.568,p=0.010$, not GPT-4o, $t(821)=1.17,p=0.24$).
Feature order in text affects generalization: In Fig. 3 we show that the order in which features are presented in text has a strong effect on how the models generalize. In general, the models are more shape-biased when shape is mentioned first, and less shape-biased if color is mentioned first (respectively; $t(1681)=6.32,p<0.001$; $t(839)=12.1,p<0.001$; $t(839)=9.23,p<0.001$). Thus, the models seem to be somewhat biased towards generalizing in accordance with the feature that is mentioned first. However, there is still a weak overall tendency to be shape biased on average.
@@ 4 Discussion
Limitations: These experiments, as a preliminary investigation, have a number of limitations. First, it would be worth evaluating the robustness of the effects across a broader range of tasks, and with other answer formats. Second, it would be worth evaluating a larger range of models, including open-source models.
@@ Acknowledgements
We thank Mike Mozer and the reviewers for helpful comments and suggestions.
@@ B.0.1 Single category
We show the results in Fig. 5. The “color match” bars show the proportion of the time the model endorses a novel probe stimulus that matches (at least one of) the exemplars on color (but does not match any on shape); the “shape match” correspondingly tests for probes that match on shape, but not color. Overall, the patterns of behavior are very noisy and inconsistent, though there is an overall tendency to favor shape matches. However, there is a hint of the expected pattern of generalization when comparing the varied-color and varied shape panels; the model tends to generalize more to matching shapes when the color of examples varies (irrespective of modality), but shows a greater degree of color generalization when the colors are consistent. However, the results when both features vary are surprisingly narrow.